# Nash-equilibrium Seeking Algorithm for Power-Allocation Games on Networks of International Relations

### v1

@@ 2.1 Notations, definitions and model setup
International relations as a signed graph: Consider a world of $n$ countries, indexed by $i\in\mathcal{V}=\{1,\dots,n\}$, whose mutual relations are fixed. For each country $i$, let $\mathcal{F}_{i}$ and $\mathcal{A}_{i}$ denote its sets of friendly and antagonistic countries, respectively. All relations are assumed bilateral and symmetric: $j\in\mathcal{F}_{i}$ if and only if $i\in\mathcal{F}_{j}$, and $j\in\mathcal{A}_{i}$ if and only if $i\in\mathcal{A}_{j}$. For convenience, we also let $i\in\mathcal{F}_{i}$ for all $i\in\mathcal{V}$. The $n$ countries together with the sets $\mathcal{F}_{1},\dots,\mathcal{F}_{n}$ and $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ form an undirected and unweighted signed graph, where nodes represent countries and positive (negative) edges represent friendly (antagonistic) relations. The graph is not assumed to be complete, i.e., $\mathcal{A}_{i}\cup\mathcal{F}_{i}=\mathcal{V}$ does NOT necessarily hold for every $i$. Throughout this paper, we use the terms “graph” and “network” (as well as “node” and “country”) interchangeably.
Power-allocation strategies: For each country $i$, let its power be a non-negative value $p_{i}$, which is assumed fixed for now. In the next section, we will model how $p_{i}$ co-evolves with the countries’ strategies. Country $i$ allocates its power either to attack its antagonistic countries or to support its friendly ones. For any $j\in\mathcal{A}_{i}$ (or $j\in\mathcal{F}_{i}$, respectively), let $x_{ij}\geq 0$ denote the amount of power that country $i$ spends on attacking (or supporting) $j$. For simplicity, both $p_{i}$ and $x_{ij}$ are assumed to take values from $\mathbb{N}$ for all $i,j\in\mathcal{V}$. This assumption of quantized power allocation does not alter the essence of the model and sometimes can even be more realistic, since national power, such as human resources or military units, is not always continuously divisible. The row vector $(x_{i1},\dots,x_{in})$ is referred to as the strategy of country $i$. Denote by $X=(x_{ij})_{n\times n}$ the strategies of all countries and call it a strategy matrix. By definition,
@@ 2.2 Existence and computation of Nash equilibrium
Now we consider the case when $X(t+1)$ is obtained from $X(t)$ via an preferable adjustment in Procedure 2 for some $i$. Since Procedure 2 requires $x_{ii}(t)=0$ and since $X(t)$ is an NSND strategy matrix, we have $p_{i}=\sum_{j\in\mathcal{A}_{i}}x_{ij}(t)=\sum_{j\in\mathcal{A}_{i}}x_{ji}(t)$. Suppose $x_{jj}(t)=0$ for any $j\in\mathcal{A}_{i}$, by Definition 4,
@@ 2.3 Extreme-case analysis: Nash equilibria in a fully antagonistic world
    If there exists a country $i\in\mathcal{V}$ such that $p_{i}>\sum_{j\neq i}p_{j}$, then there is only one safe country at any Nash Equilibrium, which is country $i$.
@@
We first prove Statement 1) by contradiction. Suppose that all countries are safe at some Nash Equilibrium $X\in\Omega$, which implies that no country could be in a more preferred situation by reallocating its power. Then, for any country $i,j\in\mathcal{V}$ with $i\neq j$, country $j$ cannot make country $i$ unsafe by assigning all $j$’s power on attacking $i$. That is,
@@
Let $j=i+1$. Here we treat node $n+1$ as node $1$ and treat node 1-1 as node $n$. By summing the left-hand side of the above inequality over all $i\in\mathcal{V}$, we obtain
@@
The proof of Statement 2) is straightforward. If there exists $i\in\mathcal{V}$ with $p_{i}>\sum_{j\neq i}p_{j}$, no matter how other countries allocate their powers, it is impossible to make country $i$ unsafe. In the meanwhile, given the power-allocation strategies of other countries, country $i$ can make all countries except itself unsafe by assigning power $p_{j}$ on attacking country $j$ for any $j\in\mathcal{V}$. Therefore, at any Nash equilibrium, there is only one safe country, which is country $i$.
@@
If the sum of all countries’ power is odd, take any three countries with positive power, denoted by $i_{1},i_{2},i_{3}$, and temporarily reduce their power by 1. Now the countries’ total power is even. We have proved that, in this case, there exists a Nash Equilibrium $X$ such that all countries are precarious. Then we return the reduced powers to the three countries and add $1$ to $x_{i_{1}i_{2}},x_{i_{2}i_{3}},and\ x_{i_{3}i_{1}}$ respectively. The new strategy matrix remains feasible. Moreover, it is still a Nash equilibrium since all countries are precarious. This completes the proof of Statement 3).

### v2

@@ 2.1 Notations, definitions and model setup
International relations as a signed graph: Consider a world of $n$ countries, indexed by $i\in\mathcal{V}=\{1,\dots,n\}$, whose mutual relations are fixed. For each country $i$, let $\mathcal{F}_{i}$ and $\mathcal{A}_{i}$ denote its sets of friendly and antagonistic countries, respectively. All relations are assumed bilateral and symmetric: $j\in\mathcal{F}_{i}$ if and only if $i\in\mathcal{F}_{j}$, and $j\in\mathcal{A}_{i}$ if and only if $i\in\mathcal{A}_{j}$. For convenience, we also let $i\in\mathcal{F}_{i}$ for all $i\in\mathcal{V}$. The $n$ countries together with the sets $\mathcal{F}_{1},\dots,\mathcal{F}_{n}$ and $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ form an undirected and unweighted signed graph, where nodes represent countries and positive (negative) edges represent friendly (antagonistic) relations. The graph is not assumed to be complete, i.e., $\mathcal{A}_{i}\cup\mathcal{F}_{i}=\mathcal{V}$ does NOT necessarily hold for every ii. Throughout this paper, we use the terms “graph” and “network” (as well as “node” and “country”) interchangeably.
Power-allocation strategies: For each country $i$, let its power be a non-negative value $p_{i}$, which is assumed fixed for now. In the next section, we will model how $p_{i}$ co-evolves with the countries’ strategies. Country $i$ allocates its power either to attack its antagonistic countries or to support its friendly ones. For any $j\in\mathcal{A}_{i}$ (or $j\in\mathcal{F}_{i}$, respectively), let $x_{ij}\geq 0$ denote the amount of power that country $i$ spends on attacking (or supporting) $j$. For simplicity, both $p_{i}$ and $x_{ij}$ are assumed to take values from $\mathbb{N}$ for all $i,j\in\mathcal{V}$. This assumption of quantized power allocation does not alter the essence of the model and sometimes can even be more realistic, since national power, such as human resources or military units, is not always continuously divisible. The row vector $(x_{i1},\dots,x_{in})$ is referred to as the strategy of country ii. Denote by $X=(x_{ij})_{n\times n}$ the strategies of all countries and call it a strategy matrix. By definition,
@@ 2.2 Existence and computation of Nash equilibrium
Now we consider the case when $X(t+1)$ is obtained from $X(t)$ via an preferable adjustment in Procedure 2 for some ii. Since Procedure 2 requires $x_{ii}(t)=0$ and since $X(t)$ is an NSND strategy matrix, we have $p_{i}=\sum_{j\in\mathcal{A}_{i}}x_{ij}(t)=\sum_{j\in\mathcal{A}_{i}}x_{ji}(t)$. Suppose $x_{jj}(t)=0$ for any $j\in\mathcal{A}_{i}$, by Definition 4,
@@ 2.3 Extreme-case analysis: Nash equilibria in a fully antagonistic world
    If there exists a country $i\in\mathcal{V}$ such that $p_{i}>\sum_{j\neq i}p_{j}$, then there is only one safe country at any Nash Equilibrium, which is country ii.
@@
We first prove Statement 1) by contradiction. Suppose that all countries are safe at some Nash Equilibrium $X\in\Omega$, which implies that no country could be in a more preferred situation by reallocating its power. Then, for any country $i,j\in\mathcal{V}$ with $i\neq j$, country $j$ cannot make country $i$ unsafe by assigning all $j$’s power on attacking ii. That is,
@@
Let $j=i+1$. Here we treat node $n+1$ as node 11 and treat node 1-1 as node $n$. By summing the left-hand side of the above inequality over all $i\in\mathcal{V}$, we obtain
@@
The proof of Statement 2) is straightforward. If there exists $i\in\mathcal{V}$ with $p_{i}>\sum_{j\neq i}p_{j}$, no matter how other countries allocate their powers, it is impossible to make country $i$ unsafe. In the meanwhile, given the power-allocation strategies of other countries, country $i$ can make all countries except itself unsafe by assigning power $p_{j}$ on attacking country $j$ for any $j\in\mathcal{V}$. Therefore, at any Nash equilibrium, there is only one safe country, which is country ii.
@@
If the sum of all countries’ power is odd, take any three countries with positive power, denoted by $i_{1},i_{2},i_{3}$, and temporarily reduce their power by 1. Now the countries’ total power is even. We have proved that, in this case, there exists a Nash Equilibrium $X$ such that all countries are precarious. Then we return the reduced powers to the three countries and add 11 to $x_{i_{1}i_{2}},x_{i_{2}i_{3}},and\ x_{i_{3}i_{1}}$ respectively. The new strategy matrix remains feasible. Moreover, it is still a Nash equilibrium since all countries are precarious. This completes the proof of Statement 3).
2.4 Inverse Inference of Unknown Edge Weights in Network Connections
In complex network systems, the strength of connections between nodes is often heterogeneous. Varying edge weights reflect the diversity of interactions among nodes. Such disparities in weighting not only influence the evolutionary dynamics of the system as a whole, but also largely determine the role and functional importance of each node within the network.
In this context, inverse inference plays an indispensable role. Since these weights often cannot be directly observed or measured, relying solely on prior information is insufficient for their accurate characterization. Therefore, it is necessary to perform inverse inference based on observed data. Through inverse inference, one can recover the hidden distribution of weights from Nash equilibrium trajectories, thereby extracting the actual intensity of interactions between nodes. More importantly, inverse inference not only reveals the true relationships among nodes but also enhances the predictive power and interpretability of the model, ensuring that research outcomes remain robust in more complex and uncertain environments.
Building upon the static power-allocation game model, we further extend our study to a more general form of influence weights between nodes. We modify the preference axiom (2) to the following form:
$\sum_{j\in E_{i}}w_{ij}I_{s_{j}(X)\leq 0}$ $+\sum_{l\in A_{i}}w_{il}I_{s_{l}(X)\geq 0}\geq$
$\sum_{j\in E_{i}}w_{ij}I_{s_{j}(Y)\leq 0}+\sum_{l\in A_{i}}w_{il}I_{s_{l}(Y)\geq 0},$
where the elements $w_{ij}$ of theweight matrix $\mathbf{W}$ denotes the importance of node $j$ to node ii. Our objective is to utilize a series of observed approximate Nash equilibrium strategy matrices from real-world data to inversely infer the differences in weights between nodes. In the following, we propose an optimization algorithm to address this problem through detailed analysis and the inspiration mainly comes from [paganGameTheoreticalInference2019a].
Suppose we have an observed matrix $\mathbf{X}^{*}$. For any node $i$, we aim to deduce the influence weights $\mathbf{w}_{i}=(w_{i1},w_{i2},\dots,w_{in})$ from other nodes on node ii. Let $\mathbf{X}=(\mathbf{x}_{i},\mathbf{x}_{-i})$, where $\mathbf{x}_{i}=(x_{i1},x_{i2},\dots,x_{in})$ represents the strategy chosen by node $i$, and $\mathbf{x}_{-i}$ denotes the strategy profile of all other nodes. Define $\mathcal{X}_{i}=\{\mathbf{x}_{i}\in\mathbb{R}^{m}\mid\sum_{j=1}^{m}x_{ij}=p_{i}\}$ as the set of all feasible strategies for node ii. The utility of node $i$ is given by:
$V_{i}(\mathbf{x}_{i},\mathbf{x}_{-i},\mathbf{w}_{i})=\sum_{j=1}^{m}w_{ij}I_{s_{i}(\mathbf{X})\geq 0}$ $+\sum_{j\in A_{i}}w_{ij}I_{s_{j}(\mathbf{X})\geq 0}$
$+\sum_{j\in E_{i}}w_{ij}I_{s_{j}(\mathbf{X})\leq 0}.$
According to variational inequality theory, when the strategy profile of all nodes constitutes a Nash equilibrium, no node $i$ can unilaterally change its own strategy to improve its utility. For a Nash equilibrium $\mathbf{X}^{*}=(\mathbf{x}_{i}^{*},\mathbf{x}_{-i}^{*})$, define:
$e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})=V_{i}(\mathbf{x}_{i},\mathbf{x}_{-i}^{*},\mathbf{w}_{i})-V_{i}(\mathbf{x}_{i}^{*},\mathbf{x}_{-i}^{*},\mathbf{w}_{i}).$
Note that $e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$ is linear with respect to each component of $\mathbf{w}_{i}$. For any feasible $\mathbf{x}_{i}$, $e_{i}$ should be non-positive. However, whenever the weight parameter $\mathbf{w}_{i}$ leads to a strategy $\mathbf{x}_{i}$ that violates the Nash equilibrium condition, the function $e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$ takes a positive value. This positive violation can be interpreted as an error, potentially arising from factors such as the bounded rationality of nodes. Therefore, we introduce an error term defined as $e_{i}^{+}(\mathbf{x}_{i},\mathbf{w}_{i})=\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\}$, representing the degree of violation of strategy $\mathbf{x}_{i}$ compared to the Nash equilibrium strategy $\mathbf{x}_{i}^{*}$ under parameter $\mathbf{w}_{i}$. A common approach to measure the distance from Nash equilibrium is to integrate the squared error term $e_{i}^{+}(\mathbf{x}_{i},\mathbf{w}_{i})$ over the entire action space, which aligns with the concept of the $L_{2}$ norm [paganGameTheoreticalInference2019a]:
$d_{i}(\mathbf{w}_{i}):=\left(\int_{\mathcal{X}_{i}}(e_{i}^{+}(\mathbf{x}_{i},\mathbf{w}_{i}))^{2}d\mathbf{x}_{i}\right)^{1/2}.$
The following lemma establishes the continuous differentiability and convexity of $d_{i}(\mathbf{w}_{i})$, providing the theoretical foundation for finding its minimum.
Lemma 2.4.
(Continuous Differentiability and Convexity of $d_{i}(\mathbf{w}_{i})$) When $\mathbf{w}_{i}$ takes values in a bounded set $\mathcal{W}$, $d_{i}(\mathbf{w}_{i}):\mathcal{W}\subseteq\mathbb{R}^{m}\to\mathbb{R}$ is a continuously differentiable convex function.
Proof 2.5.
First, assume that all components in $\mathcal{W}$ do not exceed $\hat{W}$. Note that for any $\mathbf{x}_{i}\in\mathbb{R}^{n}$, $\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\}$ has a continuous first-order derivative with respect to $\mathbf{w}_{i}$, and its gradient is:
$\nabla_{\mathbf{w}_{i}}$ $(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}$
$=\begin{cases}0,&\text{if }e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})<0,\\ 2e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\nabla_{\mathbf{w}_{i}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i}),&\text{otherwise}.\end{cases}$
This holds because $e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$ is linear in the components of $\mathbf{w}_{i}$, making $\nabla_{\mathbf{w}_{i}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$ continuous with respect to $\mathbf{w}_{i}$. Define $\mathcal{X}_{+}(\mathbf{w}_{i})=\{\mathbf{x}_{i}\in\mathcal{X}\mid e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})>0\}$ and $\mathcal{X}_{-}(\mathbf{w}_{i})=\{\mathbf{x}_{i}\in\mathcal{X}\mid e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\leq 0\}$. For any $j\in\{1,2,\dots,m\}$, since $(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}$ is integrable with respect to $\mathbf{x}_{i}$ for any fixed $w_{ij}$, and
$\left|\frac{\partial}{\partial w_{ij}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}\right|\leq 4m\hat{W}^{2},$
it follows from Theorem 2.27 in [follandRealAnalysis2013] that:
$\frac{\partial d_{i}(\mathbf{w}_{i})}{\partial w_{ij}}$ $=\int_{\mathcal{X}}\frac{\partial}{\partial w_{ij}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}d\mathbf{x}_{i}$
$=\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}\frac{\partial}{\partial w_{ij}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}d\mathbf{x}_{i}$
$\ \ \ \ +\int_{\mathcal{X}_{-}(\mathbf{w}_{i})}\frac{\partial}{\partial w_{ij}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}d\mathbf{x}_{i}$
$=\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}\frac{\partial}{\partial w_{ij}}(e_{i}(\mathbf{x}_{i},\mathbf{w}_{i}))^{2}d\mathbf{x}_{i}$
$=\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}2e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})d\mathbf{x}_{i}$
$=\int_{\mathcal{X}}2\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\}\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})d\mathbf{x}_{i}.$
Note that since $e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$ is linear in $\mathbf{w}_{i}$, $\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$ is independent of $\mathbf{w}_{i}$, implying that $\frac{\partial d_{i}(\mathbf{w}_{i})}{\partial w_{ij}}$ is continuous in $\mathbf{w}_{i}$. Therefore, the gradient in vector form is:
$\nabla_{\mathbf{w}_{i}}d_{i}(\mathbf{w}_{i})=\int_{\mathcal{X}}2\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\}\nabla_{\mathbf{w}_{i}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})d\mathbf{x}_{i}.$
Next, we compute the Hessian matrix of $d_{i}(\mathbf{w}_{i})$. Note that $\frac{\partial}{\partial w_{ij}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}$ is integrable with respect to $\mathbf{x}_{i}$ for any fixed $w_{ij}$, and for any $k\in\{1,2,\dots,m\}$, $\frac{\partial^{2}}{\partial w_{ij}\partial w_{ik}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}$ exists, with
$\left|\frac{\partial^{2}}{\partial w_{ij}\partial w_{ik}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}\right|\leq 4.$
Thus, by Theorem 2.27 in [follandRealAnalysis2013]:
$\frac{\partial^{2}d_{i}(\mathbf{w}_{i})}{\partial w_{ij}\partial w_{ik}}$
$\ \ \ \ \ =\frac{\partial}{\partial w_{ik}}\int_{\mathcal{X}}2\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\}\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})d\mathbf{x}_{i}$
$\ \ \ \ \ =\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}\frac{\partial}{\partial w_{ik}}\left[2e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\right]d\mathbf{x}_{i}$
$\ \ \ \ \ =2\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}[\frac{\partial}{\partial w_{ik}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$
$\ \ \ \ \ \ \ \ \ +e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\frac{\partial^{2}}{\partial w_{ij}\partial w_{ik}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})]d\mathbf{x}_{i}$
$\ \ \ \ \ =2\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}\frac{\partial}{\partial w_{ik}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})d\mathbf{x}_{i}.$
Therefore, the Hessian matrix is $H(d_{i}(\mathbf{w}_{i}))=2\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}\nabla_{\mathbf{w}_{i}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})(\nabla_{\mathbf{w}_{i}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i}))^{T}d\mathbf{x}_{i}$, which is positive semi-definite because for any $\mathbf{y}\in\mathbb{R}^{n}$:
$\mathbf{y}^{T}H(d_{i}(\mathbf{w}_{i}))\mathbf{y}=2\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}\left(\mathbf{y}^{T}\nabla_{\mathbf{w}_{i}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\right)^{2}d\mathbf{x}_{i}\geq 0.$
Hence, $d_{i}(\mathbf{w}_{i})$ is a convex function.
From Lemma 3, $d_{i}(\mathbf{w}_{i})$ is a continuously differentiable convex function on any bounded set. This allows us to propose the following meaningful optimization problem:
Problem 2.6.
(Minimum Nash Equilibrium Distance Problem) Given a signed network with $n$ nodes, for all nodes $i$, find the weight parameter $\hat{\mathbf{w}}_{i}$ that solves the following optimization problem:
$\min_{\mathbf{w}_{i}\in\mathcal{W}}(d(\mathbf{w}_{i}))^{2}$
$\text{s.t. }\mathbf{w}_{i}\in\mathcal{W}.$
Although the objective function is convex, the lack of guaranteed strict convexity makes direct optimization challenging. We first approximate the Riemann integral by discretizing the feasible strategy space into a set of discrete points. Suppose we select feasible strategies $\mathbf{x}_{i}^{1},\mathbf{x}_{i}^{2},\dots,\mathbf{x}_{i}^{n_{i}}$. The optimization problem then transforms into:
Problem 2.7.
(Discrete Minimum Nash Equilibrium Distance Problem) Given a signed network with $n$ nodes, for all nodes $i$, find the weight parameter $\hat{\mathbf{w}}_{i}$ that solves the following optimization problem:
$\min_{\mathbf{w}_{i}\in\mathcal{W}}\sum_{j=1}^{n_{i}}(e_{i}^{(j,+)}(\mathbf{w}_{i}))^{2}$
$\text{s.t. }\mathbf{w}_{i}\in\mathcal{W}.$
In complex network modeling and inference, if an edge weight is inferred to be non-zero, it implies a substantive interaction between the corresponding nodes. To avoid spurious connections caused by observational noise or numerical approximation, it is necessary to set a fixed positive threshold, ensuring that all non-zero weights strictly exceed this value. This constraint not only theoretically guarantees the interpretability of the edge’s existence but also effectively distinguishes genuine interactions from artifacts induced by random perturbations, thereby enhancing the scientific validity and robustness of the model results. Furthermore, the model may face overfitting issues, causing it to rely excessively on limited observed data and lose its ability to capture underlying general patterns. To address this, we introduce a regularization term in the inference process to balance model fit and complexity within the optimization objective. Consequently, we ultimately solve the following optimization problem to infer the weights influencing node $i$ via inverse inference:
Problem 2.8.
(Discrete Minimum Nash Equilibrium Distance Problem) Given a signed network with $n$ nodes, for all nodes $i$, find the weight parameter $\hat{\mathbf{w}}_{i}$ that solves the following optimization problem:
$\min_{\mathbf{w}_{i}\in\mathcal{W}}\sum_{j=1}^{n_{i}}(e_{i}^{(j,+)}(\mathbf{w}_{i}))^{2}+\|\mathbf{w}_{i}\|_{2}^{2}$
$\text{s.t. }w_{ij}\geq 0.01,\quad\forall 1\leq j\leq m.$
Note that the first part of the objective function is linear in $\mathbf{w}_{i}$, while the second part is a convex quadratic function. As the components of $\mathbf{w}_{i}$ tend to positive infinity, the objective function also tends to positive infinity, ensuring that if a minimum exists, it must lie within a bounded set. Moreover, since the objective function is strictly convex over this bounded set and the variables lie in a convex set, the solution to this optimization problem exists and is unique.